# Exact Coset Sampling forQuantum Lattice Algorithms

### v6

@@ 3 The new Step $9^{\dagger}$
1:Input: A lattice instance $L_{q}^{\perp}(\mathbf{A})$ arising from a chosen-secret LWE instance as in chen2024quantum (i.e., $\mathbf{A}=[2p_{1}\mathbf{t}\mid\mathbf{U}^{\mathsf{T}}\mid\mathbf{I}_{m}]$) together with public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
@@
3:Parameters: Re-tuned Karst-wave configuration $(c,D,r,s,t,u,P,M,Q)$ (see Remark 3.1).
@@
5:Fix absolute constants $C_{\mathrm{rank}}>1$, $C_{\mathrm{try}}\geq 1$, and $C_{\mathrm{samp}}\geq 1$.
6:Set the target sample count $N\leftarrow\lceil C_{\mathrm{rank}}(n-1)\rceil$ and the batch budget $T_{\max}\leftarrow\lceil C_{\mathrm{try}}n\rceil$.
7:Set a per-batch oracle-call cap $B_{\max}\leftarrow\lceil C_{\mathrm{samp}}\cdot N\rceil$.
8:Precompute $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$ $\triangleright$ Exists by AC2-AC3. Used inside AC4.
9:for $\mathrm{trial}=1$ to $T_{\max}$ do
10:   Initialize $\mathcal{S}\leftarrow\emptyset$ and $\mathrm{calls}\leftarrow 0$.
11:  while $|\mathcal{S}|<N$ and $\mathrm{calls}<B_{\max}$ do
12:     $\mathrm{calls}\leftarrow\mathrm{calls}+1$.
13:     $\bm{u}_{Q}\leftarrow\textsc{Step9DaggerSample}(D,M,Q,\mathrm{inv})$ $\triangleright$ Algorithm 2.
14:   if $\bm{u}_{Q}\neq\bm{0}$ then
15:        Append $\bm{u}_{Q}$ as a row of $\mathcal{S}$.
16:   end if
17:  end while
18:  if $|\mathcal{S}|<N$ then
19:     continue $\triangleright$ Sampling cap exceeded (wrong guess or atypical transcript).
20:  end if
21:   $\mathbf{b}_{Q}\leftarrow\textsc{RecoverBModQ}(\mathcal{S},Q)$ $\triangleright$ Algorithm 4.
22:  if $\mathbf{b}_{Q}=\textsc{fail}$ then
23:     continue
24:  end if
25:   $(s,e)\leftarrow\textsc{ExtractAndVerifyLWE}(\mathbf{b}_{Q},\mathbf{A},q,\beta,p_{1},\dots,p_{\kappa})$ $\triangleright$ Algorithm 3.
26:  if $(s,e)\neq\textsc{fail}$ then
27:     return $(s,e)$
28:  end if
29:end for
30:return fail
2:Oracle: One copy of $|\widetilde{\varphi}_{7}\rangle$ on coordinate registers $X\in(\mathbb{Z}_{M})^{n}$ (Definition 2.7).
3:Output: A reduced sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ (possibly $\bm{0}$).
4:Query the state-preparation oracle to obtain $|\widetilde{\varphi}_{7}\rangle$ on $X$.
5:Measure $\tau\leftarrow X_{1}\bmod D^{2}$ $\triangleright$ Deterministic conditioned on the transcript by AC1/AC4.
6:Compute $v_{1,Q}\leftarrow\operatorname{ctr}(\tau)\bmod Q$.
7:Apply the $\tau$-controlled diagonal unitary $U_{\mathrm{corr}}^{(\tau)}$ on $X_{1}$ as defined in AC4 (using $\mathrm{inv}$ and $v_{1,Q}$).
8:Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to $X$ and measure $\bm{u}\in\mathbb{Z}_{M}^{n}$.
9:return $\bm{u}_{Q}\leftarrow\bm{u}\bmod Q$.
@@
1:Input: $\mathbf{b}_{Q}\in\mathbb{Z}_{Q}^{n}$, the defining matrix $\mathbf{A}$ of $L_{q}^{\perp}(\mathbf{A})$, and public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
@@
3:Let $m\leftarrow$ the number of rows of $\mathbf{A}$ and $n\leftarrow$ the number of columns of $\mathbf{A}$.
4:Let $\ell\leftarrow n-m-1$ (so that $\mathbf{A}$ has block form $2p_{1}\mathbf{t}\mid\mathbf{U}^{\mathsf{T}}\mid\mathbf{I}_{m}]$ with $\mathbf{U}\in\mathbb{Z}_{q}^{\ell\times m}$ in the chosen-secret embedding).
@@
14:for $\eta=2$ to $\kappa$ do
15:  if $b_{\eta}\neq 2p_{1}p_{\eta}$ in $\mathbb{Z}$ then
@@
21:if $\mathbf{A}\mathbf{b}\not\equiv\bm{0}\pmod{q}$ then
@@
3:Let $U\in(\mathbb{Z}_{Q})^{N\times(n-1)}$ be columns $2..n$ of $\mathcal{S}$, and let $\bm{y}\in(\mathbb{Z}_{Q})^{N}$ be column 11 of $\mathcal{S}$.
4:Solve $U\bm{z}\equiv\bm{y}\pmod{Q}$ and test uniqueness (e.g. via Smith normal form).
@@
As in chen2024quantum, the small modulus is chosen as $Q=(c{+}1)\|\mathbf{b}\|^{2}$, where $\mathbf{b}$ is the (unknown) unique shortest vector of $L_{q}^{\perp}(\mathbf{A})$. We run a classical outer loop over all polynomially many candidates $B_{\mathrm{guess}}$ for $\|\mathbf{b}\|^{2}$. For each guess we set $Q\leftarrow(c{+}1)B_{\mathrm{guess}}$, choose an odd $D$ with $\gcd(D,Q)=1$, derive $M=2D^{2}Q$ (hence $M_{2}=D^{2}Q$), and complete the remaining Karst-wave parameters as in Definition 2.1. We then execute Algorithm 1 for this guess.
@@
where $s$ is the LWE secret and $e$ is the error vector. Because $\|\mathbf{b}\|_{\infty}\leq 2p_{1}\beta\log n<Q/2$ in the parameter regime of interest, the reduction map $\mathbf{b}\mapsto\mathbf{b}\bmod Q$ is injective on the centred cube that contains the true shortest vector. Therefore, once Algorithm 1 recovers $\mathbf{b}\bmod Q$, the canonical lift to $(-Q/2,Q/2]^{n}$ recovers $\mathbf{b}$ as an integer vector. Dividing the last n1n-1 coordinates by $2p_{1}$ (after checking divisibility) yields $s$ and $e$, and the check $\mathbf{A}\mathbf{b}\equiv 0\pmod{q}$ certifies correctness.
@@ 3.1 Correctness under Additional Conditions
Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the (unnormalized) amplitude equals
@@
so the outcome probabilities are proportional to the squared magnitude of the inner sum (the global phase $\omega_{M}^{\langle\bm{u},\bm{v}^{\ast}\rangle}$ disappears).
@@
where $t(\bm{u}):=\langle\mathbf{b},\bm{u}\rangle\bmod Q\in\mathbb{Z}_{Q}$ and we identify $t(\bm{u})$ with its representative in $\{0,\dots,Q-1\}$ when used inside the exponential. Therefore $S_{E}(\bm{u})$ factorizes as
$S_{E}(\bm{u})=\Bigl(\sum_{j\in\mathbb{Z}}\alpha_{E}(j)\,e^{2\pi ij\,t(\bm{u})/Q}\Bigr)\cdot\Bigl(\sum_{\bm{k}\in\mathcal{K}}\beta_{E}(\bm{k})\,i^{\|\bm{k}\|^{2}}\,(-1)^{\langle\bm{u},\bm{k}\rangle}\Bigr)=\widehat{\alpha}_{E}(\langle\mathbf{b},\bm{u}\rangle/Q)\cdot G_{E}(\bm{u}),$
where $G_{E}(\bm{u})$ depends only on the parity pattern of $\bm{u}_{[2..n]}$. Thus
$\Pr[\bm{u}]\ \propto\ |S_{E}(\bm{u})|^{2}=|\widehat{\alpha}_{E}(\langle\mathbf{b},\bm{u}\rangle/Q)|^{2}\cdot|G_{E}(\bm{u})|^{2}.$
@@
Because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, for every fixed $\bm{u}_{[2..n]}\in\mathbb{Z}_{M}^{n-1}$ and every $t$ there are exactly $M/Q=2D^{2}$ values of $u_{1}\in\mathbb{Z}_{M}$ such that $\bm{u}\in T_{t}$. Since $G_{E}(\bm{u})$ does not depend on $u_{1}$, the total grid mass $\sum_{\bm{u}\in T_{t}}|G_{E}(\bm{u})|^{2}$ is the same for every $t$. Hence the total probability mass of $T_{t}$ is proportional to $|\widehat{\alpha}_{E}(t/Q)|^{2}$, up to negligible truncation error.
By the leakage bound (2.1) (good transcript) and this proportionality, the total off-resonant probability mass satisfies
@@
Assume AC1-AC5 and run Algorithm 1 in the correct iteration of the outer norm-guessing loop. Each sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ appended to $\mathcal{S}$ by Algorithm 1 is non-zero and satisfies
@@
except with probability $\operatorname{negl}(n)$ over the choice of a good transcript and the internal randomness of the measurements. Solving the resulting linear system modulo $Q$ (using the known normalization $b_{1}=-1$) recovers $\mathbf{b}\bmod Q$ and hence $\mathbf{b}$ as an integer vector. The LWE secret and error are then extracted and verified in polynomial time with overwhelming probability.
@@
In each oracle call within Algorithm 1, the sampling primitive Step9DaggerSample (Algorithm 2) measures $\tau=X_{1}\bmod D^{2}$ and applies the diagonal gate $U_{\mathrm{corr}}^{(\tau)}$. Lemma 2.4 (which uses the canonical-gauge requirement in AC4) shows that this removes the quadratic chirp and produces the chirp-free state $|\varphi_{7}\rangle$ assumed in AC5.
@@
Fix such a successful batch. Because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, the projection $H\to\mathbb{Z}_{Q}^{n-1}$ given by $\bm{u}\mapsto\bm{u}_{[2..n]}$ is a bijection. Hence the accepted vectors $\bm{u}_{Q,[2..n]}$ are i.i.d. (up to a negligible statistical error from the rare off-resonant outputs) uniform over $\mathbb{Z}_{Q}^{n-1}\setminus\{\bm{0}\}$. Let $U\in\mathbb{Z}_{Q}^{N\times(n-1)}$ be the matrix with these rows and let $\bm{y}$ be the vector of first coordinates. Then the true unknown $\mathbf{b}_{[2..n]}\bmod Q$ satisfies $U\mathbf{b}_{[2..n]}\equiv\bm{y}\pmod{Q}$.
With overwhelming probability $U$ is injective as a map $(\mathbb{Z}_{Q})^{n-1}\to(\mathbb{Z}_{Q})^{N}$. Indeed, for any prime $p\mid Q$, the reduction $U\bmod p$ is (up to a negligible error from excluding the all-zero row) an i.i.d. uniform random matrix over the field $\mathbb{Z}_{p}$ with $N\geq C_{\mathrm{rank}}(n-1)$ rows, and standard rank bounds give $\Pr[\operatorname{rank}_{\mathbb{Z}_{p}}(U\bmod p)<n-1]\leq p^{-\Omega(n)}$. If $\operatorname{rank}_{\mathbb{Z}_{p}}(U\bmod p)=n-1$ and $U\bm{z}\equiv\bm{0}\pmod{p^{e}}$, then reducing modulo $p$ gives $\bm{z}\equiv\bm{0}\pmod{p}$; writing $\bm{z}=p\bm{z}_{1}$ and iterating shows $\bm{z}\equiv\bm{0}\pmod{p^{e}}$, so $\ker(U\bmod p^{e})=\{\bm{0}\}$. By the Chinese remainder theorem, $\ker(U\bmod Q)=\{\bm{0}\}$. Therefore the system has a unique solution $\bm{z}\in\mathbb{Z}_{Q}^{n-1}$, which Algorithm 4 computes (e.g. via Smith normal form) and uses to set $\mathbf{b}_{Q}=[-1,\bm{z}^{\top}]^{\top}=\mathbf{b}\bmod Q$.
Finally, the bound $\|\mathbf{b}\|_{\infty}<Q/2$ from (R1) makes the centered lift from $\mathbb{Z}_{Q}^{n}$ to $(-Q/2,Q/2]^{n}$ injective on the relevant region, so lifting $\mathbf{b}_{Q}$ recovers $\mathbf{b}$ as an integer vector. The subsequent divisibility, planted-prefix, shortness, and lattice-membership checks (Algorithm 3) succeed for the correct guess and reject incorrect guesses, as in chen2024quantum. ∎

### v7

@@ 3 The new Step $9^{\dagger}$
1:Input: A chosen-secret LWE instance $(\mathbf{U},\mathbf{t})$ as in chen2024quantum over modulus $q$ (equivalently, the lattice $L_{q}^{\perp}(\mathbf{A})$ with $\mathbf{A}=[2p_{1}\mathbf{t}\mid\mathbf{U}^{\mathsf{T}}\mid\mathbf{I}_{m}]$), together with public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
@@
3:Parameters: A re-tuned Karst-wave configuration $(c,D,r,s,t,u,P,M,Q)$ as in Definition 2.1 (chosen for the current norm guess; see Remark 3.1).
@@
5:Let $\ell\leftarrow$ the number of rows of $\mathbf{U}$ and $m\leftarrow$ the number of columns of $\mathbf{U}$; set $n\leftarrow 1+\ell+m$.
6:Fix absolute constants $C_{\mathrm{rank}}>1$, $C_{\mathrm{try}}\geq 1$, and $C_{\mathrm{samp}}\geq 1$.
7:Set the target sample count $N\leftarrow\lceil C_{\mathrm{rank}}(n-1)\rceil$ and the batch budget $T_{\max}\leftarrow\lceil C_{\mathrm{try}}n\rceil$.
8:Set a per-batch oracle-call cap $B_{\max}\leftarrow\lceil C_{\mathrm{samp}}\cdot N\rceil$.
9:Precompute $\mathrm{inv}:=(-2D^{2})^{-1}\in\mathbb{Z}_{Q}$ $\triangleright$ Exists by AC2-AC3. Used inside AC4.
10:for $\mathrm{trial}=1$ to $T_{\max}$ do
11:   Initialize $\mathcal{S}\leftarrow\emptyset$ and $\mathrm{calls}\leftarrow 0$.
12:  while $|\mathcal{S}|<N$ and $\mathrm{calls}<B_{\max}$ do
13:     $\mathrm{calls}\leftarrow\mathrm{calls}+1$.
14:     $\bm{u}_{Q}\leftarrow\textsc{Step9DaggerSample}(D,M,Q,\mathrm{inv})$ $\triangleright$ Algorithm 1.
15:   if $\bm{u}_{Q}\neq\bm{0}$ then
16:        Append $\bm{u}_{Q}$ as a row of $\mathcal{S}$.
17:   end if
18:  end while
19:  if $|\mathcal{S}|<N$ then
20:     continue $\triangleright$ Sampling cap exceeded (wrong guess or atypical transcript).
21:  end if
22:   $\mathbf{b}_{Q}\leftarrow\textsc{RecoverBModQ}(\mathcal{S},Q)$ $\triangleright$ Algorithm 4.
23:  if $\mathbf{b}_{Q}=\textsc{fail}$ then
24:     continue
25:  end if
26:   $(s,e)\leftarrow\textsc{ExtractAndVerifyLWE}(\mathbf{b}_{Q},\mathbf{U},\mathbf{t},q,\beta,p_{1},\dots,p_{\kappa})$ $\triangleright$ Algorithm 3.
27:  if $(s,e)\neq\textsc{fail}$ then
28:     return $(s,e)$
29:  end if
30:end for
31:return fail
@@
1:Input: $\mathbf{b}_{Q}\in\mathbb{Z}_{Q}^{n}$, a chosen-secret LWE instance $(\mathbf{U},\mathbf{t})$ with $\mathbf{U}\in\mathbb{Z}_{q}^{\ell\times m}$ and $\mathbf{t}\in\mathbb{Z}_{q}^{m}$, and public parameters $(q,\beta,p_{1},\dots,p_{\kappa})$.
@@
3:Let $\ell\leftarrow$ the number of rows of $\mathbf{U}$ and $m\leftarrow$ the number of columns of $\mathbf{U}$; set $n\leftarrow 1+\ell+m$.
4:(Fail if $\kappa-1>\ell$.) $\triangleright$ Chosen-secret prefix must fit in the $\ell$-dimensional secret.
@@
14:for $i=2$ to $\kappa$ do
15:  if $b_{i}\neq 2p_{1}p_{i}$ in $\mathbb{Z}$ then
@@
21:if $\mathbf{U}^{\mathsf{T}}s+e\not\equiv\mathbf{t}\pmod{q}$ then
@@
3:Let $W\in(\mathbb{Z}_{Q})^{N\times(n-1)}$ be the submatrix of $\mathcal{S}$ consisting of columns $2..n$, and let $\bm{y}\in(\mathbb{Z}_{Q})^{N}$ be column 11 of $\mathcal{S}$.
4:Solve $W\bm{z}\equiv\bm{y}\pmod{Q}$ and test uniqueness (e.g. via Smith normal form (kannan1979polynomial)).
@@
As in chen2024quantum, the small modulus is chosen as $Q=(c{+}1)\|\mathbf{b}\|^{2}$, where $\mathbf{b}$ is the unknown unique shortest vector of $L_{q}^{\perp}(\mathbf{A})$. We run a classical outer loop over all polynomially many candidates $B_{\mathrm{guess}}$ for $\|\mathbf{b}\|^{2}$. For each guess we set $Q\leftarrow(c{+}1)B_{\mathrm{guess}}$, choose an odd $D$ with $\gcd(D,Q)=1$, derive $M=2D^{2}Q$ (hence $M_{2}=D^{2}Q$), and complete the remaining Karst-wave parameters as in Definition 2.1. We then execute Algorithm 2 for this guess.
@@
[…]
@@ 3.1 Correctness under Additional Conditions
Fix $E$. Apply $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ to the chirp-free state Equation (1.2). For each $\bm{u}\in\mathbb{Z}_{M}^{n}$, the unnormalized amplitude equals
@@
so the outcome probabilities are proportional to the squared magnitude of the inner sum.
@@
where $t(\bm{u}):=\langle\mathbf{b},\bm{u}\rangle\bmod Q\in\mathbb{Z}_{Q}$ and we identify $t(\bm{u})$ with its representative in $\{0,\dots,Q-1\}$ when used inside the exponential. Equivalently, if we write $\bm{u}=\bm{u}_{Q}+Q\bm{t}$ with $\bm{u}_{Q}=\bm{u}\bmod Q$, then $\langle\bm{u},\Delta\rangle-\langle\bm{u}_{Q},\Delta\rangle=\langle Q\bm{t},2D^{2}\mathbf{b}\rangle\in M\mathbb{Z}$, so $\omega_{M}^{j\langle\bm{u},\Delta\rangle}=\omega_{M}^{j\langle\bm{u}_{Q},\Delta\rangle}$ for all integers $j$. Therefore $S_{E}(\bm{u})$ factorizes as
$S_{E}(\bm{u})=\Bigl(\sum_{j\in\mathbb{Z}}\alpha_{E}(j)\,e^{2\pi ij\,t(\bm{u})/Q}\Bigr)\cdot\Bigl(\sum_{\bm{k}\in\mathcal{K}}\beta_{E}(\bm{k})\,i^{\|\bm{k}\|^{2}}\,(-1)^{\langle\bm{u},\bm{k}\rangle}\Bigr)=\widehat{\alpha}_{E}(t(\bm{u})/Q)\cdot G_{E}(\bm{u}),$
where $G_{E}(\bm{u})$ depends only on the parity pattern of $\bm{u}_{[2..n]}$. For notational clarity, write $t(\bm{u}):=\langle\mathbf{b},\bm{u}\rangle\bmod Q\in\mathbb{Z}_{Q}$ and view $t(\bm{u})$ in $\{0,\dots,Q-1\}$ when used as an argument of $\widehat{\alpha}_{E}(\cdot)$. Then the envelope factor above is $\widehat{\alpha}_{E}(t(\bm{u})/Q)$.
Thus
$\Pr[\bm{u}]\ \propto\ |S_{E}(\bm{u})|^{2}=|\widehat{\alpha}_{E}(t(\bm{u})/Q)|^{2}\cdot|G_{E}(\bm{u})|^{2}.$
@@
Because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, for every fixed $\bm{u}_{[2..n]}\in\mathbb{Z}_{M}^{n-1}$ and every $t$ there are exactly $M/Q=2D^{2}$ values of $u_{1}\in\mathbb{Z}_{M}$ such that $\bm{u}\in T_{t}$. Since $G_{E}(\bm{u})$ does not depend on $u_{1}$, the total grid mass $\sum_{\bm{u}\in T_{t}}|G_{E}(\bm{u})|^{2}$ is the same for every $t$. Hence, the total probability mass of $T_{t}$ is proportional to $|\widehat{\alpha}_{E}(t/Q)|^{2}$, up to negligible truncation error.
By the leakage bound Equation (2.1) (good transcript) and this proportionality, the total off-resonant probability mass satisfies
@@
Assume AC1-AC5 and run Algorithm 2 in the correct iteration of the outer norm-guessing loop. Each accepted sample $\bm{u}_{Q}\in\mathbb{Z}_{Q}^{n}$ appended to $\mathcal{S}$ by Algorithm 2 is non-zero by construction and satisfies
@@
except with probability $\operatorname{negl}(n)$ over the choice of a good transcript and the internal randomness of the measurements. Solving the resulting linear system modulo $Q$ recovers $\mathbf{b}\bmod Q$ and hence $\mathbf{b}$ as an integer vector. The LWE secret and error are then extracted and verified in polynomial time with overwhelming probability.
@@
In each oracle call within Algorithm 2, the sampling primitive Step9DaggerSample (Algorithm 1) measures $\tau=X_{1}\bmod D^{2}$, obtains the run-local residue $v_{1,Q}=v_{1}^{\ast}(E)\bmod Q$ from run-local side information (AC4), and applies the diagonal gate $U_{\mathrm{corr}}^{(v_{1,Q})}$. Lemma 2.4 shows that this removes the quadratic chirp and produces the chirp-free state $|\varphi_{7}\rangle$ assumed in AC5.
@@
Fix such a successful batch. Because $b_{1}=-1$ is a unit in $\mathbb{Z}_{Q}$, the projection $H\to\mathbb{Z}_{Q}^{n-1}$ given by $\bm{u}\mapsto\bm{u}_{[2..n]}$ is a bijection. Moreover, on $H$ the only vector with $\bm{u}_{[2..n]}=\bm{0}$ is $\bm{u}=\bm{0}$ (because $u_{1}\equiv 0$ then forces $u_{1}\equiv 0$), so rejecting $\bm{u}_{Q}=\bm{0}$ is equivalent to rejecting $\bm{u}_{Q,[2..n]}=\bm{0}$. Hence, the accepted vectors $\bm{u}_{Q,[2..n]}$ are i.i.d. (up to a negligible statistical error from the rare off-resonant outputs) uniform over $\mathbb{Z}_{Q}^{n-1}\setminus\{\bm{0}\}$. Let $W\in\mathbb{Z}_{Q}^{N\times(n-1)}$ be the matrix with these rows and let $\bm{y}$ be the vector of first coordinates. Then the true unknown $\mathbf{b}_{[2..n]}\bmod Q$ satisfies $W\mathbf{b}_{[2..n]}\equiv\bm{y}\pmod{Q}$.
With overwhelming probability $W$ is injective as a map $(\mathbb{Z}_{Q})^{n-1}\to(\mathbb{Z}_{Q})^{N}$. Because $Q=\operatorname{poly}(n)$, the number of distinct prime divisors of $Q$ is at most $\log_{2}Q=O(\log n)$, so a union bound over $p\mid Q$ preserves a negligible overall failure probability. Indeed, for any prime $p\mid Q$, the reduction $W\bmod p$ is (up to a negligible error from excluding the all-zero row) an i.i.d. uniform random matrix over the field $\mathbb{Z}_{p}$ with $N\geq C_{\mathrm{rank}}(n-1)$ rows, and standard rank bounds give $\Pr[\operatorname{rank}_{\mathbb{Z}_{p}}(W\bmod p)<n-1]\leq p^{-\Omega(n)}$. If $\operatorname{rank}_{\mathbb{Z}_{p}}(W\bmod p)=n-1$ and $W\bm{z}\equiv\bm{0}\pmod{p^{e}}$, then reducing modulo $p$ gives $\bm{z}\equiv\bm{0}\pmod{p}$; writing $\bm{z}=p\bm{z}_{1}$ and iterating shows $\bm{z}\equiv\bm{0}\pmod{p^{e}}$, so $\ker(W\bmod p^{e})=\{\bm{0}\}$. By the Chinese remainder theorem, $\ker(W\bmod Q)=\{\bm{0}\}$. Therefore the system has a unique solution $\bm{z}\in\mathbb{Z}_{Q}^{n-1}$, which Algorithm 4 computes and uses to set $\mathbf{b}_{Q}=[-1,\bm{z}^{\top}]^{\top}=\mathbf{b}\bmod Q$.
Finally, the bound $\|\mathbf{b}\|_{\infty}<Q/2$ from (R1) makes the centered lift from $\mathbb{Z}_{Q}^{n}$ to $(-Q/2,Q/2]^{n}$ injective on the relevant region, so lifting $\mathbf{b}_{Q}$ recovers $\mathbf{b}$ as an integer vector. The subsequent divisibility, planted-prefix, shortness, and LWE-consistency checks (Algorithm 3) succeed for the correct guess and reject incorrect guesses, as in chen2024quantum. ∎
@@
4 Related Work
Lattice-based cryptography is rooted in worst-case hardness frameworks such as Ajtai’s construction of hard lattice instances (ajtai1996generating) and subsequent worst-case-to-average-case reductions based on Gaussian measure (micciancio2007worst). Regev introduced LWE and gave quantum worst-case reductions from lattice problems such as GapSVP and SIVP to LWE (regev2009lattices). Subsequent work clarified parameter tradeoffs and established classical hardness in a range of regimes (peikert2009public; brakerski2013classical). These reductions underpin a broad ecosystem of cryptographic constructions, including trapdoors for hard lattices (gentry2008trapdoors) and fully homomorphic encryption (gentry2009fully).
For context, the fastest known worst-case SVP algorithms are classical and typically rely on sieving and/or discrete Gaussians, starting from the sieve of Ajtai-Kumar-Sivakumar (ajtai2001sieve) and later practical refinements (nguyen2008sieve), and including the $2^{n+o(n)}$-time discrete-Gaussian algorithm of aggarwal2015solving. Discrete Gaussian sampling is also a central technical tool in lattice algorithms more broadly; efficient lattice Gaussian samplers were developed, for example, by Peikert (peikert2010efficient).
On the quantum side, lattice algorithms are often organized around preparing structured coset- or line-supported superpositions and extracting information via Fourier sampling. A canonical early connection is Regev’s reduction from unique-SVP to the dihedral hidden subgroup problem (DHSP) via coset sampling (regev2004quantum), together with subexponential-time algorithms for DHSP such as Kuperberg’s sieve (kuperberg2005subexponential). Related dihedral-coset models have since appeared in lattice settings, e.g., the extrapolated dihedral coset problem (EDCP), which is closely connected to LWE (brakerski2018learning).
On the average-case state side, Chen, Liu, and Zhandry introduced a filtering framework for decoding hidden linear structure from LWE-like quantum states and used it to obtain polynomial-time quantum algorithms for several average-case variants, including EDCP in certain regimes (chen2022quantum). Subsequent work has studied both algorithms and hardness barriers for quantum LWE-state families with structured (e.g., Gaussian) amplitudes and linear/quadratic phase terms (chen2025lwe).
Chen’s Karst-wave construction (chen2024quantum) is an instance of this Fourier-sampling template with amplitude engineering via complex-Gaussian windowing, yielding a post-processing state supported on a shifted grid line whose amplitudes include a quadratic chirp in the loop index. Chen’s published post-processing (Steps 8-9) exploits additional arithmetic structure to handle the unknown shift and chirp. The present work isolates an explicit access-model promise that suffices for exact post-processing in the LWE regime: run-local access to $v_{1,Q}:=v_{1}^{\ast}(E)\bmod Q$ (AC4), enabling single-coordinate chirp cancellation followed by a direct $\mathrm{QFT}_{\mathbb{Z}_{M}}^{\otimes n}$ without learning the full offset $\bm{v}^{\ast}(E)$.